OpenStudy (anonymous):

@563blackghost 9. Peter is trying to buy fencing for the perimeter of his garden. His garden is in the shape of a rectangle with a length of 2(x+6)feet and a width of 3.5x feet. How many feet of fencing will he need to buy? Write and simplify an expression to represent this situation? What properties did you use?

OpenStudy (anonymous):

P=2(L+W) P=2(2(x+6)+3.5x P=4(x+24+3.5x P=4x+3.5x+24 P=7.5x+24

563blackghost (563blackghost):

http://openstudy.com/study#/updates/54d91012e4b0ec5104412292

OpenStudy (anonymous):

ok